### v1

# Optimized Minimal 3D Gaussian Splatting

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful representation for real-time, high-performance rendering, enabling a wide range of applications. However, representing 3D scenes with numerous explicit Gaussian primitives imposes significant storage and memory overhead. Recent studies have shown that high-quality rendering can be achieved with a substantially reduced number of Gaussians when represented with high-precision attributes. Nevertheless, existing 3DGS compression methods still rely on a relatively large number of Gaussians, focusing primarily on attribute compression. This is because a smaller set of Gaussians becomes increasingly sensitive to lossy attribute compression, leading to severe quality degradation. Since the number of Gaussians is directly tied to computational costs, it is essential to reduce the number of Gaussians effectively rather than only optimizing storage. In this paper, we propose Optimized Minimal Gaussians representation (OMG), which significantly reduces storage while using a minimal number of primitives. First, we determine the distinct Gaussian from the near ones, minimizing redundancy without sacrificing quality. Second, we propose a compact and precise attribute representation that efficiently captures both continuity and irregularity among primitives. Additionally, we propose a sub-vector quantization technique for improved irregularity representation, maintaining fast training with a negligible codebook size. Extensive experiments demonstrate that OMG reduces storage requirements by nearly 50% compared to the previous state-of-the-art and enables 600+ FPS rendering while maintaining high rendering quality. Our source code is available at https://maincold2.github.io/omg/.
Figure 1: Our approach focuses on minimizing storage requirements while using only a minimal number of Gaussian primitives. By proposing an efficient attribute representation, including sub-vector quantization, we achieve scene representations under 5 MB with 600+ FPS rendering. We visualize qualitative examples (left) and the rate-distortion curve evaluated on the Mip-NeRF 360 dataset (right). All rendering speeds were measured on an NVIDIA RTX 3090 GPU, with values in parentheses in the left visualizations measured using an NVIDIA RTX 4090 GPU.
1 Introduction
3D Gaussian Splatting (3DGS) [26] has gained popularity for fast and photorealistic 3D scene reconstruction and rendering, offering a compelling alternative to conventional methods. By leveraging tile-based parallelism to approximate NeRF’s [37] volumetric rendering, 3DGS enables significantly accelerated rendering while maintaining high visual quality. This has facilitated a wide range of applications, such as dynamic scene reconstruction [58, 55], photorealistic avatar generation [45, 38], generative models [53, 10], and city-scale rendering [27, 46], demonstrating its versatility across various domains.
3DGS adjusts the number of Gaussian primitives during training by iteratively cloning or splitting Gaussians with high positional gradients while removing low-opacity Gaussians. However, this optimization process introduces a substantial number of redundant Gaussians (over 3 million per 360 scenes [2]), leading to excessive storage requirements and computational overhead. To address this issue, various approaches have been proposed, including pruning based on rendering loss [29, 33] or importance score [43, 12] and optimized densification strategies [36]. Notably, several methods [12, 13, 61] reduce the number of Gaussians to around 0.5 million, enabling real-time rendering even on low-capacity GPUs while preserving rendering quality.
Despite these efforts, reducing the number of Gaussians alone does not sufficiently mitigate storage overhead. Each Gaussian is parameterized by 59 learnable parameters, so even with a reduced number of primitives, storage consumption remains substantial (e.g., 133 MB in Fig. 1). To address this, many works have explored compressing Gaussian attributes by leveraging vector quantization [41, 11], neural fields [50], sorting mechanisms [39], and entropy optimization [7, 54], demonstrating considerable improvements in reducing storage consumption.
However, the aforementioned compression methods typically rely on a large number of Gaussians (over one million). This is due to two major challenges when the number of Gaussians is drastically reduced: 1) each Gaussian needs to represent a larger portion of the scene, making it more susceptible to compression loss, and 2) the increased spacing between Gaussians disrupts spatial locality, leading to higher attribute irregularity and posing challenges for entropy minimization and efficient compression. Since the number of Gaussians directly impacts computational costs, including training time and rendering speed, it is crucial to develop approaches that effectively minimize the number of Gaussians while maintaining compressibility.
In this paper, we propose Optimized Minimal Gaussian representation (OMG), an efficient compression framework that operates with a minimal number of primitives. To address the irregularity of sparse Gaussians and maximize the compressibility, we employ per-Gaussian features in a novel way. Although the reduced number of Gaussians leads to a decrease in local continuity, we can still leverage the spatial correlation associated with each Gaussian’s position. Therefore, we introduce a lightweight neural field model with negligible parameters to capture the coarse spatial feature. This feature is integrated with the per-Gaussian features to represent each attribute, as shown in Fig. 2. This approach requires fewer per-Gaussian parameters than directly learning the original attributes, enabling a more compact representation. While the proposed OMG architecture effectively represents sparse Gaussians, the use of per-Gaussian features impacts storage efficiency. To mitigate this, we introduce a sub-vector quantization (SVQ, Fig. 3(c)), which splits the input vector into multiple sub-vectors and applies vector quantization to each sub-vector. This approach alleviates the computational overhead associated with large vector quantization codebooks (Fig. 3(a)) and reduces the storage burden caused by the multiple indexing stages of residual vector quantization (Fig. 3(b)), while maintaining high-precision representation.
Finally, to retain only the minimal number of Gaussians, we introduce a novel importance metric that evaluates each Gaussian’s local distinctiveness relative to its neighbors, identifying the most informative Gaussians. This metric is used alongside existing importance scoring methods based on blending weights from training views [43, 12], further reducing the number of Gaussians while preserving scene fidelity.
Extensive experimental results demonstrate that OMG achieves a 49% reduction in storage compared to the previous state-of-the-art method [50], requiring only 4.1MB for the Mip-NeRF 360 dataset [2] while preserving comparable rendering quality. Additionally, OMG utilizes only 0.4 million Gaussians, enabling 600+ FPS rendering. These results underscore the effectiveness of OMG in both compression efficiency and computational performance, demonstrating it as a highly promising approach for 3D Gaussian Splatting representation.
2 Related Work
2.1 Neural Radiance Fields
Neural Radiance Fields (NeRF) [37] introduced a pioneering approach for novel view synthesis by leveraging volumetric rendering in conjunction with Multilayer Perceptrons (MLPs) to model continuous 3D scenes. While NeRF achieves high-quality rendering, its reliance on MLP leads to inefficiencies, particularly in terms of slow training and inference times. To overcome these limitations, the following methods [14, 31] utilized explicit voxel-based representations, enabling significantly faster training compared to traditional MLP-based NeRF models. However, these approaches still suffer from slow inference speeds and impose substantial memory requirements, posing challenges for scalability and practical deployment in large-scale environments.
Compact representation. To mitigate the memory overhead while maintaining rendering fidelity, various works have been introduced, including grid factorization [5, 4, 15, 20, 16], hash grids [40, 6], grid quantization [52, 49], and pruning-based strategies [47]. Nevertheless, achieving real-time rendering for complex, large-scale scenes remains a formidable challenge. The fundamental limitation of these approaches stems from the necessity of dense volumetric sampling, which, despite optimizations, continues to constrain training and inference speed.
2.2 3D Gaussian Splatting
Recently, 3D Gaussian Splatting (3DGS) [26] has emerged as a paradigm-shifting technique for real-time neural rendering by representing a scene with 3D Gaussian primitives. 3DGS leverages customized CUDA kernels and optimized algorithms to achieve unparalleled rendering speed while preserving high image quality. Unlike volumetric methods that require dense per-ray sampling, 3DGS projects Gaussians onto the image plane and rasterizes them tile-wise, significantly improving computational efficiency. Due to its versatility, 3DGS has become a dominant paradigm in 3D representation, leading to advancements across various domains and applications, such as mesh extraction [19, 22], simultaneous localization and mapping (SLAM) [25], dynamic scene representation [35], multi-resolution rendering [59], and further improvements in rendering quality [12]. However, 3DGS requires a substantial number of Gaussians to maintain high-quality rendering. Furthermore, each primitive is represented with multiple attributes, such as covariance matrices and spherical harmonics (SH) coefficients, requiring a large number of learnable parameters. Consequently, 3DGS demands substantial memory and storage resources, often exceeding 1GB per scene in high-resolution environments.
Figure 2: The overall architecture of our proposed OMG. OMG learns per-Gaussian geometric and appearance features, applying Sub-Vector Quantization (SVQ) to all of them. The SVQ-applied geometric attributes are used for rendering, while the space feature based on the Gaussian center position is integrated into the appearance features to define the final appearance.
Reducing the number of Primitives. To alleviate the substantial computational and memory overhead of 3DGS, numerous methods have been proposed to reduce the number of Gaussians while preserving rendering quality. Several approaches follow 3DGS by pruning low-opacity Gaussians, incorporated with opacity regularization [41], anchored Gaussians [34], or hyperparameter search [39]. An alternative approach utilized binary masking techniques [29, 7, 51, 54, 56], where pruning decisions are directly learned based on rendering loss. To optimize the binary masks, Compact-3DGS [41] initially adopted STE [3], while subsequent works [62, 33] employed Gumbel-Softmax.
Another direction focuses on importance-based metrics to identify and remove redundant Gaussians. These methods primarily leverage each Gaussian’s blending weight contribution to rendering training-view images as a measure of importance [11, 12, 43, 17, 36]. LightGaussian [11] further incorporates Gaussian volume and opacity into the importance computation, while Taming 3DGS [36] integrates multiple information, including gradients, pixel saliency, and Gaussian attributes. Building upon these advancements, we introduce a novel importance metric that incorporates color distinction among neighboring Gaussians, enabling more effective selection of essential primitives.
Attribute compression. Earlier methods employed conventional compression techniques such as scalar and vector quantization (VQ) [11, 41, 42, 39, 17, 44, 57] and entropy coding [42, 39, 29, 7, 8] to reduce storage requirements. VQ-based representations have proven highly efficient by the fact that many Gaussian attributes are redundant across a scene, allowing for compact encoding. However, a large codebook leads to substantial computational overhead, increasing training time. While residual vector quantization (R-VQ) [29] can alleviate computational costs, it introduces additional storage inefficiencies due to the need for multiple code indices.
Another line of work explored structured representations, incorporating anchor-based encoding [34, 7, 56, 32] and factorization techniques [51], integrated with grid representations. Scaffold-GS [34] first introduced an anchor-based approach, where attributes of grouped Gaussians are encoded using shared anchor features and MLP-based refinements. Building upon this, subsequent methods [7, 9, 56, 32] incorporated context modeling to further improve compression rates. While showing high compression performance, these approaches require per-view processing, involving multiple MLP forward passes, which results in significant rendering latency.
Recent efforts have utilized neural field architectures to exploit the local continuity of neighboring Gaussians. Compact-3DGS [29] encodes view-dependent color, while LocoGS [50] represents all Gaussian attributes except for view-independent color. However, unlike NeRF-based representations, where exact spatial positions are used as inputs to neural fields, mapping Gaussian center points to their corresponding attributes remains challenging. This difficulty leads to the use of large neural field models to achieve accurate reconstruction. In this work, we propose a novel approach that effectively captures both the continuity and irregularities across Gaussians, enabling a more efficient and compact attribute representation.
3 Method
Background. 3DGS represents a scene using a set of $N$ Gaussians, parameterized by their attributes: center position $p\in\mathbb{R}^{N\times 3}$, opacity $o\in[0,1]^{N}$, 3D scale $s\in\mathbb{R}_{+}^{N\times 3}$, 3D rotation represented as a quaternion $r\in\mathbb{R}^{N\times 4}$, and view-dependent color modeled using spherical harmonics (SH) coefficients $h^{(0)}\in\mathbb{R}^{N\times 3}$ (0 degree for static color), $h^{(1,2,3)}\in\mathbb{R}^{N\times 45}$ (1 to 3 degrees for view-dependent color). The covariance matrix of each Gaussian $\Sigma_{n}\in\mathbb{R}^{3\times 3}$ is determined by scale $s_{n}$ and rotation $r_{n}$ attributes.
To render an image, 3D Gaussians are projected into 2D space. Each pixel color in the image $C(\cdot)$ is then rendered through the alpha composition using colors $c_{n}$ (determined by spherical harmonics under the given view direction) and the final opacity in 2D space $\alpha_{n}(\cdot)$,
$C(x)=\sum_{k=1}^{\mathcal{N}(x)}c_{k}\alpha_{k}(x)\prod_{j=1}^{k-1}(1-\alpha_{j}(x)),$ (1)
$\alpha_{n}(x)={o_{n}}\operatorname*{exp}\left(-{1\over 2}(x-p^{\prime}_{n})^{T}{\Sigma^{\prime}}_{n}^{-1}(x-p^{\prime}_{n})\right),$ (2)
where $x$ denotes a pixel coordinate and $\Sigma^{\prime}_{n},p^{\prime}_{n}$ are the projected Gaussian covariance and center position. $\mathcal{N}(x)$ represents the number of Gaussians around $x$, where the Gaussians are depth-sorted based on the given viewing direction.
3.1 Overall Architecture
OMG is designed to accurately and efficiently represent the attributes of the minimal Gaussian primitives. Existing approaches [29, 50] have leveraged neural fields to exploit the local continuity of Gaussian attributes. However, as Gaussians become sparser, local continuity between them decreases. Unlike in dense representations, where smooth transitions between Gaussians can maintain fidelity, such transitions become less feasible in sparse settings. Especially for geometry, each Gaussian covers a larger spatial region, requiring a more specific scale and rotation to accurately capture structural details. Therefore, we retain the per-Gaussian parameterization for scale $s\in\mathbb{R}_{+}^{N\times 3}$ and rotation $r\in\mathbb{R}^{N\times 4}$ as in 3DGS.
For appearance, local continuity can still be maintained even with increased sparsity. However, unlike NeRF, where the query input is a direct spatial point, mapping Gaussian center points to corresponding appearances is inherently challenging. This requires a larger neural field model to maintain high fidelity. Conversely, entirely disregarding local continuity leads to an inefficient representation, limiting the ability to capture meaningful spatial relationships. OMG addresses these challenges by integrating per-Gaussian attributes with a neural field structure, effectively leveraging both irregularity and continuity.
To represent appearance, we learn the static feature $T\in\mathbb{R}^{N\times 3}$ and view-dependent feature $V\in\mathbb{R}^{N\times 3}$, as per-Gaussian attributes. As illustrated in Fig. 2, each feature is concatenated with the space feature $F_{n}$, derived from each Gaussian’s center position, to generate static and view-dependent color, and opacity. The space feature itself is efficiently parameterized using positional encoding and an MLP, ensuring a highly compact representation. Formally, this process can be expressed as follows:
[…]
Figure 3: Conceptual diagram of (a) vector quantization, (b) residual vector quantization, and (c) sub-vector quantization. + and $\oplus$ denote the element-wise summation and the vector concatenation.
3.2 Sub-Vector Quantization
Vector quantization (VQ) [18] has shown high efficiency for representing Gaussian attributes, capitalizing on their inherently vectorized structure and strong global coherence across an entire scene. However, to maintain high fidelity, a large codebook size is required, inevitably resulting in substantial computational overhead and increased training complexity [60] (Fig. 3(a)). To address these issues, Residual Vector Quantization (R-VQ) [60] has been used as a hierarchical quantization strategy [29], progressively refining representations while reducing the size of each individual codebook. However, as shown in Fig. 3(b), multiple code indices per attribute result in increased storage overhead, illustrating a tradeoff between reducing per-codebook complexity and increasing overall storage requirements.
To navigate this tradeoff, we propose Sub-Vector Quantization (SVQ), which partitions the attribute vector into multiple sub-vectors and applies vector quantization separately to each component (Fig. 3(c)), motivated by Product Quantization [24]. By reducing the dimensionality of each quantized unit, SVQ allows for smaller codebooks and more efficient lookups, which can balance computational cost and storage efficiency while maintaining high fidelity. We can apply SVQ to an input vector $z\in\mathbb{R}^{ML}$, where $M$ and $L$ represent the total number of sub-vectors (partitions) and the sub-vector length, respectively. Each partition $m\in\{1,...,M\}$ has an independent codebook $C^{(m)}\in\mathbb{R}^{B\times L}$, where $B$ denotes the number of codewords per codebook. The codeword selection is based on the nearest match from $C^{(m)}$, with $C^{(m)}[j]$ representing the $j$-th codeword corresponding to the $m$-th sub-vector. More formally, SVQ-applied vector $\hat{z}$ can be formulated as follows,
$\hat{z}:=q(z;M)=\text{cat}(C^{(1)}[i_{1}],C^{(2)}[i_{2}],...,C^{(M)}[i_{M}]),$ (5)
$i_{m}=\arg\min_{j}\|z_{m}-C^{(m)}[j]\|^{2}_{2},\quad m\in\{1,...,M\},$ (6)
where $q(z;M)$ denotes applying SVQ with M sub-vectors and $i_{m}\in\{1,…,B\}$ is the selected index of $m$-th sub-vector.
SVQ ensures significantly reduced computation with small codebooks compared to VQ. We apply SVQ to geometric attributes $s_{n},r_{n}$, resulting in quantized vectors $\hat{s}_{n},\hat{r}_{n}$, which are then used for 3DGS rendering. For the appearance features $T_{n},V_{n}$, we first concatenate them and apply SVQ. The resulting quantized features are then split back into two components $\hat{T}_{n},\hat{V}_{n}$, which replace $T_{n}$ and $V_{n}$ in Eqs. 3 and 4.
Although the reduced codebook size significantly decreases computational overhead compared to VQ, the process of updating both the indices and codes at every training iteration increases training time. Moreover, we observe that as training converges, the selected codebook indices remain largely unchanged. Therefore, we adopt a fine-tuning strategy in the final 1K iterations: after initializing with K-means, we freeze the indices and finetune only the codebook using the rendering loss, without introducing any additional losses. Since K-means initialization is completed within seconds due to the small codebooks, this approach adds minimal additional training time, unlike other methods that incur significant overhead.
Table 1: Quantitative results of OMG evaluated on the Mip-NeRF 360 dataset. Baseline results are sourced from the LocoGS [50] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU. We highlight the results among compression methods by coloring the best, second-best, and third-best performances.
Mip-NeRF 360
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size(MB) $\downarrow$ FPS $\uparrow$
3DGS 27.44 0.813 0.218 822.6 127
Scaffold-GS [34] 27.66 0.812 0.223 187.3 122
CompGS [41] 27.04 0.804 0.243 22.93 236
Compact-3DGS [29] 26.95 0.797 0.244 26.31 143
C3DGS [42] 27.09 0.802 0.237 29.98 134
LightGaussian [11] 26.90 0.800 0.240 53.96 244
EAGLES [17] 27.10 0.807 0.234 59.49 155
SOG [39] 27.01 0.800 0.226 43.77 134
HAC [7] 27.49 0.807 0.236 16.95 110
LocoGS-S [50] 27.04 0.806 0.232 7.90 310
LocoGS-L [50] 27.33 0.814 0.219 13.89 270
OMG-XS 27.06 0.807 0.243 4.06 350 (612)
OMG-M 27.21 0.814 0.229 5.31 298 (511)
OMG-XL 27.34 0.819 0.218 6.82 251 (416)
3.3 Local Distinctiveness for Important Scoring
OMG adopts importance scoring to identify essential Gaussians and retain a minimal number of them. Existing scoring-based pruning methods typically determine the importance of each Gaussian based on its blending weights (the values multiplied by $c_{k}$ in Eq. 1) across training-view renderings. We use two factors as our baseline metric: (1) whether it has been the most dominant contributor for at least one ray [12, 13] and (2) its total blending weight contribution across all training rays [43, 17]. Formally, we define the base importance score $\bar{I}$ as:
$\bar{I}_{i}=\begin{cases}\sum_{\rho=1}^{N^{R}}w_{i,\rho},&\text{if }\exists\rho\in\{1,...,N^{R}\}|w_{i,\rho}=\max_{j}w_{j,\rho},\\ 0,&\text{otherwise},\end{cases}$ (7)
where $w_{i,\rho}$ represents the blending weight of Gaussian $i$ for ray $\rho$ and $N^{R}$ is the number of total rays in training views.
While this score captures global importance, it does not account for redundancy among closely-positioned Gaussians. In cases where multiple Gaussians are located in close proximity, their blending weights tend to be highly similar, thus naively thresholding them can lead to two potential issues: (1) abrupt performance degradation when all similar Gaussians are simultaneously removed, and (2) redundancy when multiple Gaussians with near-identical contributions are retained.
To mitigate these issues, we propose incorporating a local distinctiveness metric into the importance computation. Specifically, we introduce an additional term that measures the appearance (static) feature similarity among neighboring Gaussians, ensuring that distinct Gaussians show high importance. The final importance score is defined as:
$I_{i}=\bar{I}_{i}\left(\frac{1}{K}\sum_{j\in\mathcal{N}_{i}^{K}}\|T_{i}-T_{j}\|_{1}\right)^{\lambda},$ (8)
where $\mathcal{N}_{i}^{K}$ denotes the set of $K$-nearest neighbors of Gaussian $i$ and $\lambda$ is a scaling factor that adjusts the sensitivity to appearance variation. As computing exact $K$-nearest neighbors for every Gaussian is computationally expensive, we approximate neighbor selection by sorting Gaussians in Morton order and selecting Gaussians with adjacent indices as their local neighbors. We remove low-importance Gaussians using CDF-based thresholding [30] with a threshold $\tau$.
Table 2: Quantitative results of OMG evaluated on the Tanks&Temples and Deep Blending datasets. Baseline results are sourced from the LocoGS [50] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU.
Tank&Temples Deep Blending
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ FPS $\uparrow$
3DGS [26] 23.67 0.844 0.179 452.4 175 29.48 0.900 0.246 692.5 134
Scaffold-GS  [34] 24.11 0.855 0.165 154.3 109 30.28 0.907 0.243 121.2 194
CompGS [41] 23.29 0.835 0.201 14.23 329 29.89 0.907 0.253 15.15 301
Compact-3DGS [29] 23.33 0.831 0.202 18.97 199 29.71 0.901 0.257 21.75 184
C3DGS [42] 23.52 0.837 0.188 18.58 166 29.53 0.899 0.254 24.96 143
LightGaussian [11] 23.32 0.829 0.204 29.94 379 29.12 0.895 0.262 45.25 287
EAGLES [17] 23.14 0.833 0.203 30.18 244 29.72 0.906 0.249 54.45 137
SOG [39] 23.54 0.833 0.188 24.42 222 29.21 0.891 0.271 19.32 224
HAC [7] 24.08 0.846 0.186 8.42 129 29.99 0.902 0.268 4.51 235
LocoGS-S [50] 23.63 0.847 0.169 6.59 333 30.06 0.904 0.249 7.64 334
OMG-M 23.52 0.842 0.189 3.22 555 (887) 29.77 0.908 0.253 4.34 524 (894)
OMG-L 23.60 0.846 0.181 3.93 478 (770) 29.88 0.910 0.247 5.21 479 (810)
4 Experiment
4.1 Implementation Details
Following the previous works, we evaluated our approach on three real-world datasets, Mip-NeRF 360 [2], Tanks&Temples [28], and Deep Blending [21]. Our model is implemented upon Mini-Splatting [12], one of the methods achieving high performance with a small number of Gaussians. We have conducted simple post-processings after training:
  • Applying 16-bit quantization to the position and compressing with G-PCC [48].
  • Huffman encoding [23] to SVQ indices.
  • Storing all the components into a single file with LZMA [1] compression.
We provide five OMG variants (XS, X, M, L, XL), adjusting storage requirements. The only factor controlling the storage is the CDF-based threshold value $\tau$ of Gaussian importance, which is set to 0.96, 0.98, 0.99, 0.999, and 0.9999 for each variant, respectively. Further implementation details are provided in the supplementary materials.
4.2 Performance Evaluation
Compression performance. Tabs. 1 and 2 compare the performance of OMG against various baseline methods on the Mip-NeRF 360, Tanks & Temples, and DeepBlending datasets. OMG consistently shows the smallest storage requirements while maintaining high performance across all datasets, achieving state-of-the-art (SOTA) results. Notably, on the Mip-NeRF 360 dataset, OMG-XS achieves nearly a 50% reduction in storage compared to the small variant of LocoGS [50], the previous SOTA compression method, while retaining PSNR and SSIM. With over 30% reduced storage, OMG-M outperforms LocoGS-S in all quality metrics. Moreover, OMG-XL surpasses LocoGS-L in all metrics, even though requiring less storage than LocoGS-S.
The qualitative results presented in Fig. 4 also demonstrate the strong performance of OMG. Despite achieving over 100× compression compared to 3DGS, OMG maintains comparable visual quality. Especially, in the bicycle scene, OMG-XS achieves over 300× compression relative to 3DGS while accurately reconstructing details that 3DGS fails to represent, resulting in a blurry area (highlighted in red) in its rendering. This superiority can be attributed to the blur split technique of our baseline model, mini-splatting [12]. Despite reducing the number of Gaussians by an additional 20% compared to mini-splatting, OMG-XS retains high visual fidelity, demonstrating its effectiveness in extreme compression scenarios.
Computational efficiency. OMG achieves remarkable efficiency alongside high performance. As shown in Tab. 3, OMG shows superior scene fidelity with significantly fewer Gaussian primitives compared to LocoGS. This reduction results in substantial rendering speed improvements of 13%, 67%, and 57% for the Mip-NeRF 360, Tank&Temples, and Deep Blending datasets (Tabs. 1 and 2), respectively, compared to LocoGS, highlighting its potential for real-time rendering on low-capacity devices. Furthermore, OMG accelerates training speed. The substantial improvement over LocoGS can be attributed to two key factors: the reduced number of Gaussians and the absence of a large neural field. By efficiently exploiting coarse spatial information through a tiny MLP, OMG achieves high computational efficiency.
Figure 4: Qualitative results of OMG compared to 3DGS. We provide per-image rendering PSNR with storage requirements for each scene.
Table 3: Efficiency comparison of OMG variants compared to LocoGS evaluated on the Mip-NeRF 360 dataset. We present training time, the number of Gaussians, and the storage requirement with rendering quality.
Method Training #Gauss Size PSNR SSIM LPIPS
LocoGS-S $\approx$1h 1.09M 7.9 27.04 0.806 0.232
LocoGS-L 1.32M 13.89 27.33 0.814 0.219
OMG-XS 20m 15s 427K 4.06 27.06 0.807 0.243
OMG-S 20m 57s 501K 4.75 27.14 0.811 0.235
OMG-M 21m 10s 563K 5.31 27.21 0.814 0.229
OMG-L 21m 32s 696K 6.52 27.28 0.818 0.220
OMG-XL 22m 26s 727K 6.82 27.34 0.819 0.218
Table 4: Ablation study of OMG using the Mip-NeRF 360 dataset. We evaluate the contribution of the space feature integration and local distinctiveness (LD) scoring.
Method PSNR SSIM LPIPS #Gauss Size
OMG-M 27.21 0.814 0.229 0.56M 5.31
w/o Space feature 26.96 0.811 0.232 0.59M 5.58
w/o LD scoring 27.09 0.813 0.230 0.57M 5.36
w/o Both 26.81 0.809 0.234 0.59M 5.59
OMG-XS 27.06 0.807 0.243 0.43M 4.06
w/o Space feature 26.85 0.804 0.246 0.44M 4.17
w/o LD scoring 26.83 0.804 0.246 0.43M 4.12
w/o Both 26.52 0.798 0.252 0.45M 4.24
Table 5: Ablation study on SVQ using the Mip-NeRF 360 dataset. We substitute SVQ to VQ.
Method Training #Gauss Size PSNR SSIM LPIPS
OMG-XS 20m 15s 427K 4.06 27.06 0.807 0.243
SVQ $\rightarrow$ VQ 21m 22s 426K 3.99 26.97 0.805 0.245
4.3 Ablation Study
OMG architecture. OMG leverages a highly compact neural field to capture coarse spatial information while reducing the number of learnable parameters per Gaussian. Tab. 4 validates the contribution of this space feature. Although the total number of Gaussians slightly increases, performance significantly degrades. The absence of spatial information introduces instability in attribute learning, hindering effective importance scoring. This trend is consistently observed in both our small and medium models, highlighting the effectiveness of the space feature despite its minimal parameter overhead.
Local distinctiveness scoring. OMG improves Gaussian pruning by incorporating local distinctiveness (LD) scoring into the importance estimation. Tab. 3 evaluates the impact of LD scoring, showing that despite a similar number of Gaussians, the inclusion of LD scoring leads to a significant performance improvement. This effect becomes even more pronounced when the target Gaussian number is lower, demonstrating that LD scoring provides an effective approach for further reducing a sparse set of Gaussians. Furthermore, when both the space feature and LD scoring are removed, the model experiences the most substantial performance drop. This indicates that the two contributions are orthogonal, independently contributing to model efficiency and performance.
Sub-vector quantization. To assess the impact of SVQ, we replaced it with VQ while maintaining the proposed clever training strategy, where K-means is performed only once before the final 1K iterations, after which only the codebook is updated. Despite this adjustment, VQ leads to a 5% increase in training time, entirely due to K-means initialization. This is because large codebooks are inevitably required for accurate representation (here, we use $2^{14}$ codes for scale and rotation and $2^{13}$ codes for each appearance feature). Nevertheless, VQ results in a decline in rendering quality, as shown in Tab. 5. To improve rendering quality, VQ requires an enlarged codebook, which is constrained by significant computational and memory overhead. Moreover, the increase in codebook size reduces storage efficiency when the number of Gaussians is small. In contrast, SVQ offers scalability of representation by flexibly adjusting the bit budget with sub-vector length in regard to the capacity-performance trade-off.
5 Conclusion
In this paper, we proposed Optimized Minimal Gaussians (OMG), a novel framework that significantly reduces the number of Gaussian primitives while maximizing compressibility and maintaining high rendering quality. By effectively identifying and preserving locally distinct Gaussians, OMG minimizes the redundancy of Gaussians with minimal loss of visual fidelity. Furthermore, our compact and precise attribute representation, combined with sub-vector quantization, enables efficient exploitation of both continuity and irregularity, ensuring high efficiency. Experimental results demonstrate that OMG reduces storage requirements by nearly 50% compared to the previous state-of-the-art method while allowing over 600 FPS rendering performance. OMG sets a new benchmark for highly efficient 3D scene representations, paving the way for future advancements in real-time rendering on resource-constrained devices.
Appendix
6 Implementation Details
All experiments were conducted using an NVIDIA RTX 4090. Our method was implemented within the mini-splatting [12] framework and trained for 30K iterations. At the 20K iteration simplification process, local distinctiveness scoring was incorporated. Scale and rotation were trained from the initial training, while appearance features were introduced at 15K iterations. At this stage, the static features were initialized using the spherical harmonics DC coefficients trained until 15K iterations, whereas view-dependent features were initialized as zero vectors.
From 29K iterations (last 1K iterations), SVQ (Sub-Vector Quantization) was applied to per-Gaussian features. As mentioned in the paper, to enhance training efficiency, K-means clustering was performed once. The assigned indices based on K-means were fixed, and only the codebooks were optimized for the remaining 1K iterations. For SVQ, different bit allocations were assigned.
  • Scale: length 1, $2^{6}$ codes for each sub-vector
  • Rotation: length 2, $2^{9}$ codes for each sub-vector
  • Appearance features: length 2, $2^{10}$ codes for each sub-vector
The length 1 SVQ applied to scale can be interpreted as scalar quantization, dynamically learning the quantization range with the codebooks. All codes in the codebook are stored with 16-FP precision.
This SVQ configuration was commonly applied across all variants from XS to XL. The model storage for each variant was determined only by the importance score threshold $\tau$, which is used for simplification at the 20K iteration, set to 0.96, 0.98, 0.99, 0.999, and 0.999, respectively.
Table 6: Ablation study on the post-processing methods applied in OMG.
G-PCC Huffman Size (MB)
- - 5.82
- 4.30
- 5.58
OMG-XS 4.06
- - 6.83
- 5.04
- 6.54
OMG-S 4.75
- - 7.66
- 5.64
- 7.33
OMG-M 5.31
- - 9.47
- 6.92
- 9.08
OMG-L 6.52
- - 9.89
- 7.25
- 9.46
OMG-XL 6.82
7 Effect of Post-Processings
As mentioned in the main paper, we applied the following two post-processing methods:
  • Compressing the 16-bit quantized position with G-PCC [48].
  • Huffman encoding [23] to SVQ indices and compressing the results with LZMA [1].
Both methods are applied losslessly, and we report the resulting storage changes in Tab. 6. When applied independently, G-PCC and Huffman encoding consistently reduce the total storage by 26-27% and 4-5% across all storage budgets, respectively. Applying both methods together also results in the overall storage reduction remaining consistent at approximately 30-32%.
Table 7: The average storage allocation for each component across OMG variants. ‘Actual size’ refers to the total size of a single file containing all components.
Attribute XS S M L XL
Position 0.93 1.08 1.20 1.43 1.52
Scale 0.83 0.97 1.09 1.33 1.41
Rotation 0.87 1.02 1.15 1.40 1.49
Appearance 1.39 1.63 1.82 2.22 2.35
MLPs 0.03 0.03 0.03 0.03 0.03
Total 4.04 4.73 5.29 6.42 6.80
Actual size 4.06 4.75 5.31 6.52 6.82
8 Storage Analysis
We conducted experiments to analyze the storage requirements of OMG for representing each attribute, as shown in Tab. 7. Across all variants, OMG allocates approximately 20-25% of the total storage to position, scale, and rotation, while around 35% is dedicated to representing appearance attributes, including static and view-dependent color as well as opacity. The four MLPs for representing local continuity and aggregating appearance attributes exhibit negligible storage requirements, even without extra compression.
9 Per-scene Results
We report per-scene results in Tab. 8 (Mip-NeRF 360 [2]) and Tab. 9 (T&T [28] and DB [21]).
Table 8: Per-scene results evaluated on the Mip-NeRF 360 [2] dataset.
Method Metric bicycle bonsai counter flowers garden kitchen room stump treehill Avg.
OMG-XS PSNR 24.95 30.90 28.40 21.32 26.42 30.81 31.09 27.00 22.60 27.06
SSIM 0.743 0.932 0.899 0.596 0.818 0.919 0.918 0.788 0.647 0.807
LPIPS 0.276 0.202 0.206 0.368 0.190 0.137 0.208 0.247 0.357 0.243
Train 18:03 20:30 24:44 19:18 18:02 23:45 20:30 17:49 19:40 20:15
#Gauss 480772 263892 310056 543034 607254 356752 281236 523821 479520 427371
Size 4.61 2.53 2.95 5.24 5.65 3.33 2.67 4.95 4.64 4.06
FPS 682 648 433 616 615 498 648 708 658 612
OMG-S PSNR 25.08 31.05 28.56 21.18 26.56 30.89 31.20 27.08 22.64 27.14
SSIM 0.750 0.936 0.903 0.602 0.826 0.921 0.922 0.792 0.650 0.811
LPIPS 0.264 0.195 0.199 0.358 0.177 0.132 0.201 0.239 0.347 0.235
Train 19:01 21:09 25:19 20:13 18:41 24:12 21:38 18:29 19:55 20:57
#Gauss 573126 310096 360930 633607 691441 412126 338884 619734 573425 501485
Size 5.46 2.94 3.41 6.10 6.43 3.83 3.19 5.83 5.54 4.75
FPS 601 585 401 555 556 462 620 601 588 552
OMG-M PSNR 25.14 31.06 28.62 21.40 26.71 31.05 31.30 27.06 22.55 27.21
SSIM 0.756 0.938 0.905 0.606 0.832 0.923 0.923 0.794 0.652 0.814
LPIPS 0.256 0.190 0.195 0.351 0.169 0.129 0.198 0.233 0.339 0.229
Train 18:58 21:01 25:44 20:35 18:51 24:18 22:14 18:31 20:22 21:10
#Gauss 646191 350999 400442 708074 772338 454908 375520 704907 649157 562504
Size 6.15 3.33 3.76 6.79 7.18 4.21 3.53 6.61 6.24 5.31
FPS 562 536 371 510 522 440 566 566 525 511
OMG-L PSNR 25.24 31.47 28.66 21.45 26.83 31.03 31.26 27.05 22.57 27.28
SSIM 0.762 0.941 0.907 0.613 0.837 0.924 0.926 0.795 0.653 0.818
LPIPS 0.241 0.183 0.189 0.338 0.160 0.126 0.191 0.226 0.329 0.220
Train 19:25 21:16 26:06 20:50 19:14 24:20 22:05 19:22 21:14 21:32
#Gauss 813561 463285 480133 859963 909961 524457 524457 869388 819435 696071
Size 7.69 4.32 4.48 8.23 8.42 4.82 4.82 8.14 7.81 6.52
FPS 476 492 332 422 422 405 539 468 414 441
OMG-XL PSNR 25.22 31.51 28.78 21.52 26.93 31.15 31.25 27.00 22.69 27.34
SSIM 0.764 0.942 0.908 0.614 0.839 0.925 0.926 0.796 0.655 0.819
LPIPS 0.239 0.182 0.187 0.334 0.157 0.126 0.191 0.224 0.324 0.218
Train 20:43 21:54 26:21 22:09 20:23 24:56 22:37 20:22 22:33 22:26
#Gauss 864124 450246 507473 922061 953050 547636 493754 920589 885229 727129
Size 8.15 4.22 4.72 8.81 8.82 5.02 4.58 8.59 8.44 6.82
FPS 430 465 324 379 422 397 512 435 384 416
Table 9: Per-scene results evaluated on the Tank&Temples [28] and Deep Blending [21] datasets.
Method Metric Tank&Temples Deep Blending
Train Truck Avg. drjohnson Playroom Avg.
OMG-M PSNR 21.78 25.25 23.52 29.37 30.18 29.77
SSIM 0.806 0.878 0.842 0.905 0.910 0.908
LPIPS 0.233 0.144 0.189 0.253 0.253 0.253
Train 12:12 11:30 11:51 17:18 14:51 16:05
#Gauss 303187 257649 330418 520385 404237 462311
Size 2.95 3.49 3.22 4.87 3.82 4.34
FPS 861 913 887 829 959 894
OMG-L PSNR 21.85 25.36 23.60 29.44 30.32 29.88
SSIM 0.811 0.881 0.846 0.907 0.912 0.910
LPIPS 0.225 0.136 0.181 0.247 0.247 0.247
Train 12:12 11:39 11:56 17:39 14:58 16:19
#Gauss 369440 442359 405900 627868 485329 556599
Size 3.58 4.28 3.93 5.86 4.55 5.21
FPS 760 780 770 745 874 810